OpenStudy (anonymous):

will medal Circle A has a diameter of 7 inches, a circumference of 21.98 inches, and an area of 38.465 square inches. The diameter of circle B is 6 inches, the circumference is 18.84 inches, and the area is 28.26 square inches. Part A: Using the formula for circumference, solve for the value of pi for each circle. (4 points) Part B: Use the formula for area and solve for the value of pi for each circle. (4 points) Part C: What observation can you make about the value of pi for circles A and B? (2 points)

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

I can help

OpenStudy (anonymous):

Circumference = (pi)*(diameter) Circumference of Circle A = 21.98 and Diameter of Circle A = 7 so (pi) = (Circumference) divided by (Diameter) = 21.98 / 7 = 3.14 which is pi same with circle b

OpenStudy (anonymous):

Thats for part A

OpenStudy (anonymous):

ok im back

OpenStudy (anonymous):

so um... what do i write for my paragraph or answer

syed98 (syedmohammed98):

lemme tag someone who can help u .. @dan815